OpenStudy (anonymous):

If y = 5 + integral (2, 2x) of e^2 dt then dy/dx = ?

OpenStudy (anonymous):

\[y=5+\int\limits_{2}^{2x}e ^{-2}dt\]

OpenStudy (anonymous):

This could be rewritten as: \[y=5+e ^{-2}\int\limits\limits_{2}^{2x}dt\]

OpenStudy (anonymous):

Oh, right! Because e is a coefficient.

OpenStudy (anonymous):

\[y = 5 + e^{-2}[t]_{2}^{2x}\]

OpenStudy (anonymous):

you know the rest right

OpenStudy (anonymous):

Oh, so you simply take the antiderivative of dt?

OpenStudy (anonymous):

let me finish the answer so it gets clearer for you \[y = 5 + e^{-2}[2x-2]=5+2e^{-2}[x-1]\]

OpenStudy (anonymous):

Hmm I didn't realize you could manipulate the dt in the integral - thank you very much!

OpenStudy (anonymous):

you are more than welcome!